OpenStudy (anonymous):

Choose the correct simplification of the expression d3 • d5. d d15 d2 d8

jimthompson5910 (jim_thompson5910):

\[d^3\cdot d^5 = d^{3+5}=d^8\] So \[d^3\cdot d^5=d^8\], which means the answer is choice D